iPad 10:01 PM Done Multiple Choices for CHEM 241-2017Fall.docx Copy Select Al Look Up Share... ovs rule. b) It needs reducing agents for boron c) It should have an anti-coplanar orientation d) It has a concerted mechanism 2. What is the element of unsaturation of CH BrNO? a) 2 b) 3) d) 5 3. What is the name of a) (E) 2-Bromo-1-phenylpropene b) (Z) 2-Bromo-l-phenylpropene c) (E) 2-Bromo-1-cyclohexylpropene d) (Z) 2-Bromo-1-cyclohexylpropene Br 4. What is the element of unsaturation of the following molecule? a) 3b4 c5 d)6 e) 7 H3C CHEM 241-41 & 51 NAME TEST #3 Oct. 31, 2017 Wednesday Score(100) 5. Which reaction of alkene would give the alcohol as its product? a) Acid-cat. hydration c) Bromination b) Alkoxymercuration-demercuration d) Free radical reaction with HBr 6. Which would be least stable alkene?
Answer #1

2) element (degree )of unsaturation is given by the formula

DU = C -H/2 -X/2 + N/2 +1

= 6-6/2 -1/2 +1/2 +1

= 4

3) option a

The name is (E) -2-bromo-1-phenylpropene

As both the higher priority groups, Br and phenyl are on opposite sides , it is E.

4) option b

While counting the unsaturation from structure,

each ring = one unsaturation

Thus DU = 2(for two rings ) + 2 double bonds

= 4

5) option a

ONly acid-catlysed hydration of alkene gives an alcohol.

The others

alkoxymercuration-demercuration gives an ether

bromination gives a dibromide

raddical addition of HBr gives a bromide

from alkene.